# P3Nav: A Unified Framework for Embodied Navigation Integrating Perception, Planning, and Prediction

### v1

@@ Abstract
In language-guided visual navigation, agents locate target objects in unseen environments using natural language instructions. For reliable navigation in unfamiliar scenes, agents must possess strong perception, planning, and prediction capabilities. Additionally, when agents revisit previously explored areas during long-term navigation, they may retain irrelevant and redundant historical perceptions, leading to suboptimal results. In this work, we introduce P3Nav, a unified framework that integrates Perception, Planning, and Prediction capabilities through Multitask Collaboration on navigation and embodied question answering (EQA) tasks, thereby enhancing navigation performance. Furthermore, P3Nav employs an Adaptive 3D-aware History Sampling strategy to effectively and efficiently utilize historical observations. By leveraging the large language models (LLM), P3Nav comprehends diverse commands and complex visual scenes, resulting in appropriate navigation actions. P3Nav achieves a 75% success rate in object goal navigation on the $\mathrm{CHORES}$-$\mathbb{S}$ benchmark, setting a new state-of-the-art performance.
@@ 1 Introduction
Embodied navigation [1], particularly language-guided visual navigation [15, 16, 22, 23, 3, 11, 6, 21] is essential for enabling robots to perform various tasks, presenting a significant challenge for researchers. Language-guided visual navigation requires agents to understand natural language instructions and autonomously navigate in unseen visual environments to locate target objects. To reliably navigate in unfamiliar visual environments, agents must effectively perceive surrounding scenes, plan strategies for reaching their goals, and predict appropriate navigation actions.
@@
Recently, significant advancements in language-guided visual navigation have been driven by the robust understanding and generalization capabilities of visual-language models (VLMs) that seamlessly integrate visual and textual information. For example, in Object Goal Navigation (ObjectNav) [22, 23, 3, 27, 33, 8, 11, 6, 21, 4, 5, 35], numerous approaches [27, 33, 8, 4, 5, 35] leveraging VLMs have achieved impressive results. These methods utilize VLMs to effectively interpret and align object categories with visual inputs, enabling agents to navigate toward specified objects with enhanced accuracy and efficiency. By combining visual and textual information, these approaches significantly enhance comprehension of complex environments and adaptation to new scenarios. While current navigation models excel at locating target objects, they often struggle to provide efficient path planning and to explain the reasoning behind their path choices, as shown in the first row of Fig. 1. This challenge is largely due to traditional navigation datasets focusing primarily on recording coordinate trajectories and neglecting the hierarchical reasoning process employed by humans. For instance, humans typically first identify rooms, then locate furniture, and finally search for specific objects.
@@
Additionally, we have observed a significant amount of redundancy in observations during navigation, as agents frequently revisit the same locations multiple times. As illustrated at the top of Fig. 2, the white circle marks areas that the agent revisits multiple times. At (1) $T=71$ and (2) $T=82$, the agent focuses on the countertop area and two chairs. At (3) $T=90$, the scene remains largely unchanged, capturing an environment similar to that at (1) and (2) once again. These observations demonstrate that the scenes within the revisit areas are highly similar, highlighting the redundancy present in the agent’s perception. Therefore, when agents revisit the same area, previous observations may become redundant and should be discarded.
To address these limitations, we present P3Nav, a unified framework for embodied navigation that integrates Perception, Planning, and Prediction. As shown in Fig. 1, we develop a Multitask Collaboration strategy that enables agents to leverage perception and planning to improve their prediction capabilities by conducting joint training on both navigation and embodied question answering (EQA) tasks. To achieve this, we have devised a novel method for constructing EQA datasets that explicitly model the decision-making process involved in navigation, with more details available in Sec. 3.2.
Furthermore, P3Nav employs an Adaptive 3D-aware History Sampling strategy to effectively and efficiently utilize historical observations in navigation. Specifically, this method selects RGB frames that do not overlap in spatial positions to serve as valid observations and introduces position-enhanced historical features that utilize the agent’s positions to augment historical semantic features with trajectory information, thereby preventing redundant exploration of the same location. As illustrated at the bottom of Fig. 2, we transform the dense and disorganized trajectory into sparse and organized steps, minimizing redundancy while preserving crucial observations at key locations. Finally, by leveraging the large language model (LLM), P3Nav can understand diverse instructions and complex visual scenes, resulting in effective navigation. Through extensive experimentation, P3Nav has demonstrated remarkable capabilities, achieving an impressive 75% success rate in ObjectNav on the $\mathrm{CHORES}$-$\mathbb{S}$ benchmark [10]. This performance not only sets a new state-of-the-art but also represents a substantial 18% absolute improvement over previous methods.
@@
Figure 2: Top: During long-term navigation, agents may revisit the same areas multiple times, making previous observations redundant. For instance, at $T=71$, $T=82$, and $T=90$, the scenes observed are very similar, highlighting the redundancy in the agent’s perception. Bottom: Using the Adaptive 3D-aware History Sampling strategy, we transform the dense and disorganized trajectory into sparse and organized steps, reducing observational redundancy while preserving visual information at key locations. The trajectory is produced by the SPOC-driven agent [10].
The main contributions of this paper include: (1) We introduce P3Nav, a unified framework integrating Perception, Planning, and Prediction, enhancing navigation by multitask collaboration on navigation and EQA tasks. (2) P3Nav employs the adaptive 3D-aware history sampling strategy, effectively utilizing historical observations by selecting non-overlapping RGB frames to reduce redundancy. (3) P3Nav achieves an 75% success rate in ObjectNav on the $\mathrm{CHORES}$-$\mathbb{S}$ benchmark, setting a new state-of-the-art, with a 18% absolute improvement over previous methods.
@@
Figure 3: Overview of P3Nav architecture. The current frame $I_{t}$ is initially processed through 2D and 3D feature extraction using the visual encoder. Then, historical features are filtered through the Adaptive 3D-aware History Sampling strategy. The visual features obtained, along with the accompanying linguistic instructions, are then fed into the large language model (LLM). Leveraging the Multitask Collaboration strategy, which significantly enhances navigation capabilities through joint training on both navigation and EQA tasks, the LLM produces two outputs via multimodal fusion: executable navigation actions and natural language answers.
@@ 2 Related Work
Embodied navigation [1] involves guiding agents through unseen environments based on human instructions. While the field includes various navigation paradigms, our research specifically focuses on multimodal tasks that combine visual perception with language processing. In this paper, we emphasize areas like ObjectNav [22] and EQA [7].
@@ 2.1 Object Goal Navigation
The research landscape of ObjectNav [22] can be systematically organized based on the scope of vocabulary and adaptability to real-world scenarios, classifying them into closed-world and open-world categories. In the closed-world category, traditional ObjectNav tasks [18, 23, 3, 27, 33, 8, 11, 6, 21, 4, 5, 35] focus on locating objects specified by predefined categorical terms, such as “chair” or “refrigerator”. Methodologies in this area primarily include end-to-end approaches [18, 23, 3, 27, 33, 8], which map sensory inputs directly to action policies, and modular architectures [11, 6, 21, 4, 5, 35], which separate perception, mapping, and control into distinct subsystems. While these strategies are effective in controlled environments, they are limited by their reliance on a fixed set of object classes. To address these limitations, Open Vocabulary Object Goal Navigation (OVON) [29] was introduced for open-world scenarios. OVON [29] enables navigation by allowing users to describe any object using natural language, such as “find a television,” thus accommodating the diverse and dynamic nature of real-world applications. Additionally, Multi-Object Navigation (MultiON) [25] extends the task from single to multiple objects, requiring agents to navigate to a sequence of targets based on language instructions. Although current ObjectNav models are effective at finding target objects, they frequently have difficulty explaining why they chose certain paths.
@@ 2.2 Embodied Question Answering
EQA [7] is an innovative task where agents explore 3D environments to answer visually grounded questions, as introduced by Das et al. [7]. This multimodal challenge requires a combination of visual perception, environment exploration, and linguistic understanding. Agents must navigate physical spaces while interpreting questions, such as ”What color is the car?” The EQA task was further developed by the MT-EQA dataset [30], which introduced scenarios involving multiple objects and locations, thereby increasing its complexity. Furthermore, the MP3D-EQA dataset [26] advanced the field by incorporating photorealistic environments with point cloud perception, challenging agents to interact with detailed 3D spaces. Recently, the HM-EQA dataset [20] refined exploration strategies with an “explore until confident” approach, promoting efficient data gathering for accurate decision-making. These aforementioned developments collectively enhanced the capabilities of agents in EQA tasks. Building on the EQA datasets, some studies [34, 32] advance the field by unifying navigation and EQA tasks into a comprehensive model.
@@ 3 Methodology
In this section, we present P3Nav, a novel and unified framework for embodied navigation that seamlessly integrates the essential components of Perception, Planning, and Prediction. First, we provide an overview of P3Nav, highlighting its central component, the Adaptive 3D-aware History Sampling strategy, as detailed in Sec. 3.1. Then, in Sec. 3.2, we introduce an innovative method to extend navigation datasets with EQA pairs that explicitly model the decision-making processes involved in navigation. Finally, we present a Multitask Collaboration strategy in Sec. 3.3, which bolsters navigation capabilities by enabling joint training on both navigation and EQA tasks.
@@ 3.1 Model
As shown in Fig. 3, P3Net integrates three specialized components for explainable robotic navigation: a visual encoder module for current observation encoding, an adaptive 3D-aware history sampling strategy for historical context modeling, and a LLM for predicting actions and answers.
@@ 3.1.1 Visual Encoder
Our visual encoder integrates 2D and 3D features to represent the current observation frame, with a dedicated emphasis on real-time perceptual cues. It consists of two specialized components: a ViT [19] for 2D feature extraction and UVFormer [17] for 3D spatial encoding.
@@
Here, the UVFormer [17] module processes three inputs: image features $X_{t}$, camera parameters $Cam$, and learnable queries $Q$ that encode spatial positions and semantic features within the robot’s 3D workspace (See [17] for architectural details). The output $UV_{t}\in\mathbb{R}^{n_{\text{UV}}\times c}$ aggregates multi-view visual information through a structured 3D feature volume, where $n_{\text{UV}}$ denotes the number of 3D visual tokens. See the supplementary material for more details.
@@ 3.1.2 Adaptive 3D-aware History Sampling
Utilizing all observed RGB frames can substantially mitigate the issue of catastrophic knowledge forgetting, which is essential for long-term navigation. However, due to GPU memory constraints and high computational costs, storing all image features extracted from these frames is often impractical. Additionally, we have observed a significant amount of redundancy in observations during navigation, as the agent frequently revisits the same locations multiple times. To address this issue, we propose a method for selecting those frames that do not overlap in spatial positions to serve as historical features, thereby reducing observational redundancy. Furthermore, we introduce position-enhanced historical features that utilize the agent’s positions to augment historical semantic features with trajectory information, preventing redundant exploration of the same location.
@@
0:  Current step $t$, Window size $W$, Threshold $\epsilon$
@@
1:  $\mathbf{V}\leftarrow\emptyset$, $\mathbf{P}\leftarrow\emptyset$
2:  $\mathbf{G}\leftarrow\text{GetAllObs}()$
@@
8:  $\mathbf{p}^{\text{ref}}\leftarrow\mathbf{G}[-1].\mathbf{p}$
9:  $k\leftarrow 0$
@@
11:  for $i\leftarrow|\mathbf{G}|-1$ to 00 do
12:     $\mathbf{p}_{i}^{\text{rel}}\leftarrow\mathbf{G}[i].\mathbf{p}-\mathbf{p}^{\text{ref}}$
13:     if $i<|\mathbf{G}|-1$ and $\exists\mathbf{q}\in\mathbf{P}:\|\mathbf{p}_{i}^{\text{rel}}-\mathbf{q}\|_{2}<\epsilon$ then
@@
17:     $\mathbf{V}\leftarrow\mathbf{V}\oplus\mathbf{MaxPool}(\mathbf{G}[i].\mathbf{v})$
@@
Specifically, the adaptive 3D-aware history sampling strategy samples image features in reverse chronological order, using the spatial position of the current frame as a reference. It retains image features with relative distances beyond a defined threshold and discards those within the threshold, thus effectively minimizing redundancy and preserving essential temporal context. As shown in Algorithm 1, our adaptive 3D-aware history sampling strategy operates through three sequential phases: initialization, 3D-aware sampling, and adaptive padding.
@@
(1) Initialization. The initialization phase (Lines 1-6) sets up the operational parameters by accepting the input parameters (current step $t$, window size $W$, threshold $\epsilon$) and initializing empty buffers for historical features $\mathbf{V}$ and relative positions $\mathbf{P}$, respectively. The GetAllObs() function returns all historical observations in $\mathbf{G}$ using a First-in-First-out (FIFO) queue, where $\mathbf{G}[0]$ corresponds to the initial time and $\mathbf{G}[-1]$ corresponds to the current time. At time step $i$, $\mathbf{G}[i]$ includes historical features $\mathbf{G}[i].\mathbf{v}$ paired with their corresponding agent absolute positions $\mathbf{G}[i].\mathbf{p}=(x,y,z)$.
@@
(2) 3D-Aware Sampling. The core 3D-aware sampling phase (Lines 8-23) processes all historical observations in $\mathbf{G}$ in reverse chronological order, from newest to oldest. First, it uses the robot’s absolute position $\mathbf{p}^{\text{ref}}$ at the current step $t$ as the spatial reference to compute relative coordinates $\mathbf{p}_{i}^{\text{rel}}$ of historical features $\mathbf{G}[i].\mathbf{v}$. Then, using the threshold $\epsilon$, it dynamically filters out spatially redundant features by removing those with distances below the threshold and preserving those above to maintain critical temporal information, as described by the condition $\|\mathbf{p}_{i}^{\text{rel}}-\mathbf{q}\|_{2}<\epsilon$. By adjusting the size of $\epsilon$, we can control the sampling density. Smaller values of $\epsilon$ result in denser sampling, allowing more continuous historical information to be retained. Concurrently, historical features $\mathbf{G}[i].\mathbf{v}$ undergo dimensionality reduction via a $\mathbf{MaxPool}$ operator before being stored. This phase terminates when the collected samples reach the target window size $W$ or $\mathbf{G}$ is empty.
@@
(3) Adaptive Padding. The adaptive padding phase (Lines 24-28) ensures fixed-length outputs by replicating the last valid entries whenever the historical features buffer $\mathbf{V}$ does not reach the designated window size $W$. Through cyclic duplication of $\mathbf{V}[-1]$ and $\mathbf{P}[-1]$, it maintains temporal coherence while fulfilling model input requirements.
Position-enhanced historical features.
@@
For obtained historical features $\mathbf{V}$, positional information is critical for recording the agent’s historical trajectory and planning efficient paths for future exploration. Hence, we obtain position-enhanced historical features $\mathbf{V_{p}}$ by using the agents’ relative positions $\mathbf{P}$ as positional encodings for the historical features $\mathbf{V}$. This integration enriches spatial semantics with 3D trajectory information, ensuring observation consistency and effectively preventing redundant exploration of previously visited locations. As agents move on the ground, their positions $\mathbf{p}$ involve constantly changing $x$ and $y$ coordinates, while the $z$ coordinate, representing height, remains constant. Consequently, we focus on encoding the agent’s positions along the $x$ and $y$ axes to effectively capture their 3D trajectory. Inspired by the work of [24], we propose a 2D encoding through axis-separated frequency projections:
@@
where $\oplus$ operation merges x/y-axis encodings into a unified spatial. For each element $\mathbf{v}$ in the historical features $\mathbf{V}$, we calculate its position encoding $PE(x,y)$. By adding these features together and passing the result through a fully connected layer, we obtain the position-enhanced historical features $\mathbf{V_{p}}\in\mathbb{R}^{n_{\text{his}}\times c}$ by enriching the historical features with positional information, where $n_{\text{his}}$ denotes the number of historical features.

### v4

# RoboTron-Nav: A Unified Framework for Embodied Navigation Integrating Perception, Planning, and Prediction

@@ Abstract
In language-guided visual navigation, agents locate target objects in unseen environments using natural language instructions. For reliable navigation in unfamiliar scenes, agents should possess strong perception, planning, and prediction capabilities. Additionally, when agents revisit previously explored areas during long-term navigation, they may retain irrelevant and redundant historical perceptions, leading to suboptimal results. In this work, we propose RoboTron-Nav, a unified framework that integrates perception, planning, and prediction capabilities through multitask collaborations on navigation and embodied question answering tasks, thereby enhancing navigation performances. Furthermore, RoboTron-Nav employs an adaptive 3D-aware history sampling strategy to effectively and efficiently utilize historical observations. By leveraging large language model, RoboTron-Nav comprehends diverse commands and complex visual scenes, resulting in appropriate navigation actions. RoboTron-Nav achieves an 81.1% success rate in object goal navigation on the $\mathrm{CHORES}$-$\mathbb{S}$ benchmark, setting a new state-of-the-art performance.
$\dagger$ Equal contribution. $\ddagger$ Corresponding author.
@@ 1 Introduction
Embodied navigation [1] is a key component of embodied artificial intelligence [12, 40, 32, 39, 19, 20], especially language-guided visual navigation [17, 18, 26, 27, 2, 11, 5, 25]. Unlike autonomous driving [13], language-guided visual navigation requires agents to comprehend natural language instructions and autonomously explore unseen visual environments to locate target objects, which presents unique challenges. Specifically, to reliably navigate in unfamiliar visual environments, agents must effectively perceive surrounding scenes, plan strategies to achieve their goals, and predict appropriate navigation actions.
@@
Recently, significant advancements in language-guided visual navigation have been driven by the robust understanding and generalization capabilities of visual-language models (VLMs) that seamlessly integrate visual and textual information. For example, in Object Goal Navigation (ObjectNav) [26, 27, 2, 31, 37, 7, 11, 5, 25, 3, 4, 41], numerous approaches [31, 37, 7, 3, 4, 41] have achieved impressive results by using VLMs to align and integrate textual and visual inputs, enabling agents to better understand complex environments and navigate to target objects more effectively. Although current navigation models excel at locating target objects, they often struggle to provide efficient path planning and explain the reasoning behind their path choices, as shown in the first row of Fig. 1. This challenge is largely due to traditional navigation datasets focusing primarily on recording coordinate trajectories and neglecting the hierarchical reasoning process employed by humans. For instance, humans typically first identify rooms, then locate furniture, and finally search for specific objects.
@@
Additionally, we have observed a significant amount of redundancy in observations during navigation, as agents frequently revisit the same locations multiple times. As illustrated at the top of Fig. 2, the white circle marks areas that the agent revisits multiple times. At (1) $T=71$ and (2) $T=82$, the agent focuses on the countertop area and two chairs. At (3) $T=90$, the scene remains largely unchanged, capturing an environment similar to that at (1) and (2) once again. These observations demonstrate that the scenes within the revisit areas are highly similar, indicating the redundancy present in the agent’s perception. Therefore, when agents revisit the same area, previous observations may become redundant and should be discarded.
@@
Figure 2: Top: During long-term navigation, agents may revisit the same areas multiple times, making previous observations redundant. For instance, at $T=71$, $T=82$, and $T=90$, the scenes observed are very similar, highlighting the redundancy in the agent’s perception. Bottom: By applying the Adaptive 3D-aware History Sampling strategy, we sample the dense and disorganized trajectory into sparse and organized steps, reducing observational redundancy while preserving visual information at key locations. The trajectory is produced by the SPOC-driven agent [10].
This scenario raises two key questions driving our research: (1) How can we clarify agents’ decision-making processes? (2) How can we enhance long-term navigation efficiency? To tackle these challenges, we propose RoboTron-Nav, a unified framework for embodied navigation that integrates perception, planning, and prediction. As illustrated in Fig. 1, RoboTron-Nav adopts a Multitask Collaboration strategy, leveraging joint training on both navigation and embodied question answering (EQA) tasks to fully harness perception and planning for improved prediction capabilities. Within this framework, the EQA task seamlessly integrates perception and planning by requiring agents to answer questions based on their historical observations and navigational trajectories. To facilitate this, we construct EQA datasets that explicitly model the decision-making processes involved in navigation, empowering agents to engage in deep reasoning and task planning akin to human cognition, and ultimately enhancing navigation interpretability. More details about our novel EQA dataset construction method are provided in Sec. 3.2.
Furthermore, RoboTron-Nav employs an Adaptive 3D-aware History Sampling strategy to effectively and efficiently minimize observation redundancy by controlling the density and diversity of historical sampling in both spatial and semantic domains. Specifically, we consider the RGB frames as valid observations only if they are either spatially non-adjacent to other frames, or spatially adjacent but captured from different viewpoints. In addition, RoboTron-Nav introduces position-enhanced historical features that utilize the agent’s positions to augment historical semantic features with trajectory information, thereby preventing redundant exploration of the same location. As illustrated at the bottom of Fig. 2, we sample the dense and disorganized trajectory to obtain sparse and organized steps, minimizing redundancy while preserving crucial observations at key locations. Finally, by leveraging the large language model (LLM), RoboTron-Nav can understand diverse instructions and complex visual scenes, resulting in effective navigation. Through extensive experimentation, RoboTron-Nav has demonstrated remarkable capabilities, achieving an impressive 81.1% success rate (SR) in ObjectNav on the $\mathrm{CHORES}$-$\mathbb{S}$ benchmark [10]. This performance not only sets a new state-of-the-art but also represents a substantial 9% absolute improvement over previous methods.
To summarize, we make the following contributions:
  • We introduce RoboTron-Nav, a unified framework integrating perception, planning, and prediction, enhancing navigation by multitask collaboration on navigation and EQA tasks.
  • RoboTron-Nav utilizes an adaptive 3D-aware history sampling strategy, which minimizes observational redundancy by regulating the density and diversity of historical sampling across both spatial and semantic domains.
  • RoboTron-Nav achieves an 81.1% SR in ObjectNav on the $\mathrm{CHORES}$-$\mathbb{S}$ benchmark, establishing a new state-of-the-art, with a 9% absolute improvement over previous methods.
@@
Figure 3: Overview of RoboTron-Nav architecture. The current frame $I_{t}$ is initially processed through 2D and 3D feature extraction using the visual encoder. Then, historical features are filtered through the Adaptive 3D-aware History Sampling strategy. The visual features obtained, along with the accompanying linguistic instructions, are then fed into the large language model (LLM). Leveraging the Multitask Collaboration strategy, which significantly enhances navigation capabilities through joint training on both navigation and EQA tasks, the LLM produces two outputs via multimodal fusion: executable navigation actions and natural language answers.
@@ 2.1 Object Goal Navigation
The ObjectNav [26] research landscape can be categorized into closed-world and open-world settings. Closed-world ObjectNav [21, 27, 2, 31, 37, 7, 11, 5, 25, 3, 4, 41] focuses on locating objects from a fixed set of object categories (e.g., “chair”). Approaches include end-to-end methods [21, 27, 2, 31, 37, 7] that map sensory inputs directly to actions, and modular architectures [11, 5, 25, 3, 4, 41] that separate perception, mapping, and control. While effective in controlled settings, these methods are limited by their reliance on predefined object classes. To address this limitation, Open Vocabulary Object Goal Navigation (OVON) [33] allows users to specify any object using natural language (e.g., “find a television”), enabling greater adaptability to real-world scenarios. Although current ObjectNav models are effective at finding target objects, they often struggle to explain their navigation decisions.
@@ 2.2 Embodied Question Answering
EQA [6, 13] tasks require agents to explore 3D environments and answer visually grounded questions, integrating visual perception, exploration, and language understanding. For example, agents need to navigate physical spaces to answer questions such as “What color is the car?” The EQA task developed by the MT-EQA dataset [34] introduces scenarios involving multiple objects and locations, thereby increasing its complexity. Furthermore, the MP3D-EQA dataset [30] advances the field by incorporating photorealistic environments with point cloud perception, challenging agents to interact with detailed 3D spaces. Recently, the HM-EQA dataset [24] refines exploration strategies with an “explore until confident” approach, promoting efficient data gathering for accurate decision-making. These aforementioned developments collectively enhance the capabilities of agents in EQA tasks. Based on EQA datasets, some studies [38, 36] advance the field by unifying navigation and EQA tasks into a comprehensive model.
@@ 3 Methodology
In this section, we present RoboTron-Nav, a unified framework for embodied navigation that integrates perception, planning, and prediction. We first overview RoboTron-Nav and its core adaptive 3D-aware history sampling strategy (Sec. 3.1). Next, we introduce a method to augment navigation datasets with EQA pairs for explicit decision modeling (Sec. 3.2). Finally, we describe a multitask collaboration strategy (Sec. 3.3) that enhances navigation by jointly training on navigation and EQA tasks.
@@ 3.1 Model
As shown in Fig. 3, RoboTron-Nav integrates three components: a visual encoder for current observations, adaptive 3D-aware history sampling for context modeling, and an LLM for action and answer prediction.
@@ 3.1.1 Visual Encoder
Our visual encoder integrates 2D and 3D features to represent the current observation frame, with a dedicated emphasis on real-time perceptual cues. It consists of two specialized components: a ViT [22] for 2D feature extraction, and a UVFormer [19] for 3D spatial encoding.
@@
Here, the UVFormer [19] module processes three inputs: image features $X_{t}$, camera parameters $Cam$, and learnable queries $Q$ that encode spatial positions and semantic features within the robot’s 3D workspace. The output $UV_{t}\in\mathbb{R}^{n_{\text{UV}}\times c}$ aggregates multi-view visual information through a structured 3D feature volume, where $n_{\text{UV}}$ denotes the number of 3D visual tokens. See the supplementary material for more details.
@@ 3.1.2 Adaptive 3D-aware History Sampling
Utilizing all observed RGB frames can substantially mitigate the issue of catastrophic knowledge forgetting, which is essential for long-term navigation. However, due to GPU memory constraints and high computational costs, storing all image features extracted from these frames is often impractical. Additionally, we have observed a significant amount of redundancy in observations during navigation, as the agent frequently revisits the same locations multiple times. To address this issue, we propose an adaptive 3D-aware history sampling strategy, which selects sparse and organized features based on historical position and viewpoint information, thereby reducing observational redundancy. As shown in Algorithm 1, we sample historical observations by using relative distance to determine spatial adjacency and semantic similarity to identify similar viewpoints. Our adaptive 3D-aware history sampling strategy operates through three sequential phases: initialization, 3D-aware sampling, and adaptive padding.
@@
0: Current step $t$, Window size $W$, Relative Position Threshold $\epsilon$, Semantic Similarity Threshold $\tau$
@@
[…]
@@
[…]
@@
11:for $i\leftarrow|\mathbf{G}|-1$ to 0 do
12:  ${p}_{i}^{\text{rel}}\leftarrow\mathbf{G}[i].{p}-{p}^{\text{ref}}$
13:  if $i<|\mathbf{G}|-1$ and $\exists\,j\in[0,k]:\ \|{p}_{i}^{\text{rel}}-\mathbf{P}[j]\|_{2}<\epsilon\;$ and $\cos\left(\mathbf{G}[i].{v},\,\mathbf{V}[j]\right)>\tau$ then
@@
17:  $\mathbf{V}\leftarrow\mathbf{V}\oplus\mathbf{MaxPool}(\mathbf{G}[i].{v})$
@@
Initialization. The initialization phase (Lines 1-6) sets up the operational parameters by accepting the input parameters (current step $t$, window size $W$, relative position threshold $\epsilon$, and semantic similarity threshold $\tau$) and initializing empty buffers for historical features $\mathbf{V}$ and relative positions $\mathbf{P}$, respectively. The GetAllObs function returns all historical observations (i.e., $X_{0},\ldots,X_{t-1}$ in Fig. 3), which are stored in $\mathbf{G}$ as a First-in-First-out (FIFO) queue, where $\mathbf{G}[0]$ corresponds to the initial time and $\mathbf{G}[-1]$ to the current time. At time step $i$, $\mathbf{G}[i]$ includes historical features $\mathbf{G}[i].{v}$ paired with their corresponding agent absolute positions $\mathbf{G}[i].{p}=(x,y,z)$.
@@
3D-Aware Sampling. The core 3D-aware sampling phase (Lines 8-23) processes all historical observations in $\mathbf{G}$ in reverse chronological order, from newest to oldest. First, it uses the robot’s absolute position ${p}^{\text{ref}}$ at the current step $t$ as the spatial reference to compute relative coordinates ${p}_{i}^{\text{rel}}$ of historical features $\mathbf{G}[i].{v}$. Then, using the relative position threshold $\epsilon$ and the semantic similarity threshold $\tau$, the method dynamically filters out redundant visual observation. Specifically, for each candidate feature at time step $i$, we compare it with all previously selected historical features at time step $j$: if $|{p}_{i}^{\text{rel}}-\mathbf{P}[j]|<\epsilon$ (i.e., adjacent in spatial position) and $\cos\left(\mathbf{G}[i].{v},\mathbf{V}[j]\right)>\tau$ (i.e., captured from similar viewpoints), feature $i$ is excluded from the historical set. By adjusting the value of $\epsilon$ and $\tau$, we can control the density and diversity of historical sampling in the spatial and semantic domains, respectively. Concurrently, historical features $\mathbf{G}[i].{v}$ undergo dimensionality reduction via a $\mathbf{MaxPool}$ operator before being stored to $\mathbf{V}$. This phase terminates when the collected samples reach the target window size $W$ or $\mathbf{G}$ is empty.
@@
Adaptive Padding. The adaptive padding phase (Lines 24-28) ensures fixed-length outputs by replicating the last valid entries whenever the historical features buffer $\mathbf{V}$ does not reach the designated window size $W$. Through cyclic duplication of $\mathbf{V}[-1]$ and $\mathbf{P}[-1]$, it maintains temporal coherence while fulfilling model input requirements.
@@
Afterwards, we introduce position-enhanced historical features that utilize the agent’s positions to augment historical semantic features with trajectory information, preventing redundant exploration of the same location. For obtained historical features $\mathbf{V}$, positional information is critical for recording the agent’s historical trajectory and planning efficient paths for future exploration. Hence, we obtain position-enhanced historical features by using the agents’ relative positions $\mathbf{P}$ as positional encodings for the historical features $\mathbf{V}$. This integration enriches spatial semantics with 3D trajectory information, ensuring observation consistency and effectively preventing redundant exploration of previously visited locations. As agents move on the ground, their positions p{p} involve constantly changing $x$ and $y$ coordinates, while the $z$ coordinate, representing height, remains constant. Consequently, we focus on encoding the agent’s positions along the $x$ and $y$ axes to effectively capture their 3D trajectory. Inspired by [29], we propose a 2D encoding through axis-separated frequency projections:
@@
where $\oplus$ operation merges xx/$y$-axis encodings into a unified spatial. We add the positional encoding $PE(x,y)$ to each historical feature in $\mathbf{V}$, and then pass these features through a fully connected layer to obtain the position-enhanced historical features ${V_{p}}\in\mathbb{R}^{n_{\text{his}}\times c}$ , where $n_{\text{his}}$ denotes the number of historical features.